A room contains air at 20 C and a total pressure of 95 kPa with a relative humidity of 40 percent. The gas constant for air is 0.287 kJ/(kg-K). The mass of dry air in the room is 30 kg. What is the volume of the room in m3 :
Answer #1

P T= 20°C (ov) 293k, = 956pa, Ⓡ=40% 011004 Given total Maryair B0kg , R=0-287 KS/- From table: for T= 20c we have had ComposWe have Pengan V = math 94.069x V= 3010-2874293 İN= 26-8vm?